Instrument Log 

 

GRAMS OT††††† C1 Optical Trailer††Lamont, OK

 

Date

Technician

Comments

10/3/00

JT

Per direction received in the SST Weekly Conference call the GRAMS instruments are to be shut down. Instrument was shut down on 10/3/2000. We are awaiting deinstallation directions from the Mentor.

7/28/00

KT

At 1428 GMT the Run / Shut down toggle switch was found in the shut down mode. At 1430 GMT the Run mode was selected. There was no response from the computer. At 1435 GMT rebooted the computer. At 1439 GMT the computer came on line and the shadow band had started rotating. RTS at 1443 GMT.

7/24/00

KT

At 1735 GMT removed the East instrument cover to check PV and SV readings. The PV reading was reading 96. While further inspection indicated that if the controller box was moved from side to side the PV value kept on changing. After discussing with Chris Martin the situation the controller box was removed and the controller box that was removed from the GRAMS RCF was installed in itís place. After controller box was installed the PV value was observed at a steady 83. RTS at 1840 GMT.

7/13/2000

CAM

GRAMS OT shadow ring controller board is needed for troubleshooting problems with GRAMS RCF. Removed GRAMS OT enclosure covers at 1515 GMT. Powered the radiometer electronics off and removed the shadow ring controller board. Returned the board after troubleshooting GRAMS RCF. Restored radiometer electronics power and RTS at 1545 GMT.

6/15/2000

CAM

GRAMS OT TDDR was removed on 6/6/00 due to a failed shadow ring motor. A replacement motor was received from Ken Black and needed to be installed. Installed the new motor and tested OK in the lab. Shut GRAMS OT down at 1902 GMT. Removed the enclosure covers and installed the TDDR. Connected the power and data cables and powered GRAMS on at 1925 GMT. The shadow ring was observed to be operational but the system shut down shortly thereafter due high temperature inside the enclosure with the covers off. Installed the enclosure covers and RTS at 1945 GMT. The following data collection was successful at 2002 GMT. RTS status O.K.

6/6/2000

CAM

Daily status check found that the TDDR shadow band was not rotating. Powered the system down at 1810 GMT. Removed the entire TDDR/shadow band assembly to the lab. Troubleshooting found that the 27VDC motor had failed. There are no spares available at this time. Installed the cover plate over the TDDR mounting hole and system RTS at 2025 GMT with the TDDR removed. Will procure a replacement motor.

4/20/00

KT

The pump breaker was tripped. At 1815 GMT reset the breaker. RTS at 1827 GMT.

4/18/2000

CAM

The new vacuum pump system has been installed and needs to be tested before activation. Tested the GRAMS27 software and found that it runs OK in normal data acquisition mode. Tested the pump system by activating the manual pump down sequence. The sequence passed through the first two steps of switching the vacuum gauge on and opening valve 1, but encountered a turbo pump fault as step three (switch on turbo controller) was activated. Troubleshooting found that the turbo controller is emitting a constant 5 volt fault signal which doesn't allow the pump to come on and stops the sequence. Ken Black removed the turbo controller and will return it to Sandia for further testing. Disabled the new GRAMS software and pump system and GRAMS C1 RTS running the old GRAMS25 software at 1800 GMT.

4/17/2000

Mentor

Ken Black and John Smith on site to install new GRAMS vacuum pump system. System to be off line most of the day. Also created new GRAMS26 and GRAMS27 directories so Scott Faas can FTP the new GRAMS software to the system. GRAMS26 will correct an error in the calculation of sunrise and sunset during daylight savings time. GRAMS27 contains the time calculation fix plus implements software for the automatic control of the vacuum pump system.

4/12/00

KT

At 1842 GMT placed grams OT pump back in service. RTS at 1917 GMT.

4/4/2000

CAM

SDS comlog reports indicated that GRAMS OT had not successfully collected since 0401 GMT on 4/1 due to an FTP problem. Attempted to monitor GRAMS OT remotely using SMS software but the system could not be reached. Removed the GRAMS cover and found the following message on the computer: An exception OE has occurred at 0028:C001B942 in VxD IFSMGR(01) + 000038FE. This was called from 0028:C1023725 in VxD VTDI(01) + 00000185. It may be possible to continue normally, press any key to continue. Attempted to shut the computer down using the start menu several times, but each time an error message indicated that Explorer had performed an illegal operation and would be shut down. Finally had to power the computer off from the desktop at 1400 GMT. Restarted the computer and system RTS status O.K. at 1405 GMT.

3/31/2000

KT

At 1855 GMT the grams pump was removed for repairs. RTS at 1925 GMT.

3/31/2000

CAM

GRAMS C1 radiometer electronics were powered off at 1524 GMT to remove the shadow band controller board to aid in T/S problems with GRAMS RCF shadow band. Reinstalled the shadow band controller board in GRAMS C1 after GRAMS RCF T/S completed. GRAMS C1 RTS status O.K. at 1700 GMT. There was no failure with the GRAMS C1 unit.

3/29/00

GRS

The SDS indicates that it has lost communications with the instrument. Troubleshooting of the instrument found that the instrument seemed to be running O.K. with the exception of the TDDR. The TDDR shadow band was not turning. Shut down the program at 2148 GMT. Power cycled the system at 2150. The system booted normally at 2153. The shadow band started and the data looks reasonable. RTS 2155 GMT

3/7/2000

CAM

GRAMS OT TDDR shadow band was not rotating properly. The band was stopped in the top center (vertical) position. The band would move slightly at times, but never completed the full arc. Removed the enclosure back cover and found that the set screw which secures the shadow band sprocket to the rotating shaft had once again come loose. The Allen head set screw was worn out and could not be tightened sufficiently. Removed the Allen head set screw and replaced with a Phillips head screw. Properly aligned the sprocket on the shaft and secured tightly with the new screw. Replaced the covers and RTS status O.K. at 1625 GMT.

3/1/2000

CAM

GRAMS OT TDDR shadow band was not rotating. Checked the instrument status remotely using SMS and found that the system was operational. Removed the enclosure back cover and found that the sprocket that turns the chain to move the shadow band was slipping on the shaft. The set screw which secures the sprocket to the shaft was loose. This allowed the shaft to turn freely inside the sprocket without ever moving it. Disconnected the motor power, aligned the shadow band at the correct position and tightened the set screw as much as possible. The head of the set screw is nearly stripped and the wrench slips when the screw becomes tight. Reconnected power and the shadow band returned operational at 2105 GMT. Will continue to monitor status and try to obtain a replacement set screw. Replaced cover and RTS status O.K. at 2110 GMT.

2/15/00

CAM

The FSBR and TSBR were pulled and returned to Sandia for repair on 2/8/00. The radiometers were not operating properly at that time. John Smith arrived on site on 2/15/00 with the repaired FSBR and TSBR radiometers after internal problems had been corrected. Removed the GRAMS enclosure lids and powered the radiometer electronics off at 1630 GMT. Installed the FSBR and TSBR radiometers and restored power. Monitored all channels via the GRAMS25 software. All radiometers are now fully operational. Replaced the enclosure covers and RTS status O.K. at 1700 GMT.

2/8/00

CAM

Ken Black returned and installed the FSBR, TSBR, and TDDR radiometers after having them modified for the vacuum system. The IOTech unit was also reinstalled after having been tested in GRAMS RCF for a day. GRAMS C1 system was powered on after all components had been reinstalled. Monitoring the GRAMS25 software revealed that the FSBR TTL lights were not changing and the value was stuck at 6.666E-6. Also, the TSBR temperature maxed out and stayed constant at 4095 after a couple of minutes. A smoke smell could also be detected coming from the GRAMS enclosure. Powered everything down and checked for damaged components but could not locate anything burned. Checked the IOTech relay card and FSBR chopper motor relay and found them to be operational. Powered the radiometer electronics off and swapped the FSBR and TSBR cables at the radiometers. Now the FSBR was connected to the TSBR cabling and electronics and vice versa. Restored radiometer power and found that the observed problems remained with the associated radiometer. In this configuration the TSBR TTL lights were not changing and it's value was stuck at 6.666E-6 on the monitor. The FSBR TTL lights were changing, however the temperature now read constant 4095. Since the heads were actually swapped, this suggests a problem with the radiometers. Ken Black removed the FSBR and TSBR back to Sandia for further analysis. GRAMS C1 RTS at 2100 GMT with only the TDDR in place.

2/7/2000

CAM

GRAMS RCF radiometers were reinstalled on 2/2 after being modified by mentor. At that time it was observed that the system program would not run. The IOTech Daqbook 100 A/D interface was suspected to be the problem. Removed the IOTech unit from GRAMS RCF on 2/7. Inspection of the IOTech revealed that the Daqbook 100 motherboard and one of the DBK21 digital I/O expansion boards had sustained major water damage. It appears that the GRAMS enclosure had leaked water during the time that the radiometers were removed for calibration. Instrument mentor requested that Site Ops remove the IOTech from the GRAMS OT and install it in GRAMS RCF to confirm operation of GRAMS RCF and the newly modified radiometers. (GRAMS OT radiometers are out for modifications at this time and the system is off line). Removed GRAMS OT IOTech and installed in GRAMS RCF and powered the system on at 2000 GMT. GRAMS RCF immediately began to operate normally. Mentor requested that GRAMS RCF be left to operate in this manner overnight. Ken Black will be on site 2/8 to reinstall the GRAMS OT radiometers, and at that time the GRAMS OT IOTech will be returned to its original location and GRAMS OT will be brought back on line. Initiated a PR for new motherboard and digital I/O board to replace the failed units.

2/2/00

Ken Black

System shut down at 1930 GMT and all sensors removed for calibration. System off line at this time.

1/21/00

KT

At 2141 GMT the psi valves on the pump were adjusted due to the pump psi value reading 1. After adjustments were made psi value stayed at 6 psi. RTS at 2148 GMT.

1/14/00

KT

The pump value was reading 2psi. At 1854 GMT powered off pump. At 1856 GMT powered on pump. Value still hadnít changed. RTS at 1859 GMT.

1/7/00

KT

At 1544 GMT observed that the pump was reading 3 psi. After inspection of the pump the value stayed the same. Troubleshooting still in progress. RTS at 1546 GMT.

12/30/99

KT

At 1511 GMT during weekly inspection of the pump the value is still reading 3. After inspection, and adjustment of the valve the pump reading never changed. RTS at 1520 GMT.

12/23/99

KT

During weekly inspection the pump value is reading 3 psi. After inspection value still the same but pump is still operating. RTS at 1725 GMT.

10/27/99

KT

Due to high winds the filter and filter cover were blown out and were found laying next to the instrument. Reinstalled filter and filter cover. RTS at 1746 GMT.

10/25/99

KT

At 1839 GMT found filter and cover on the ground next to the instrument. Cleaned filter and reinstalled filter and cover. RTS at 1843 GMT.

10/1/8/99

CAM

SDS comlog reports indicate that no files have been collected from GRAMS C1 since 0501 GMT on the 16th. The error message indicates Collection failed. No files to collect. Removed the GRAMS enclosure cover at 1540 GMT and found that the PC was running, but the LabView program appeared to be hung up. Rebooted the PC and attempted to install SMS software for remote monitoring capabilities. Was unable to see the GRAMS PC across the network using SMS. Rebooted the computer again and RTS without SMS at 1640 GMT. The program started running normally.

8/25/99

CAM

Daily status check found that the TDDR shadow band was not making its full range of movement. The band was essentially vertical and moving back and forth only a few degrees. The band was not moving enough to entirely cover the TDDR sensors surface. Removed the GRAMS enclosure cover and observed that the sprocket which drives the chain to move the shadow band was slipping on the shaft. The sprocket was only catching on the shaft occasionally. Removed the TDDR assembly to the Optical Trailer. Tightened the two Allen set screws which secure the sprocket to the shaft. Reinstalled the TDDR assembly and checked shadow band operation. The bands arc of rotation was not centered about the sensor. The band was traveling too far to the east and was not moving far enough west to completely unshade the sensor. Disconnected the motor power when the shadow band was at its farthest point west. Loosened the sprocket set screws and manually rotated the shadow band to center its rotation arc about the sensor. Tightened the sprocket set screws and connected the motor power. Verified correct shadow band rotation and replaced the enclosure cover. RTS status O.K. at 1700 GMT.

8/24/99

CAM

Daily status check on 8/19 found that the TDDR shadow band was not rotating. T/S at that time found that the shaft between the motor assembly and shadow band gears had come loose at the motor end. Attempts to reconnect the shaft were unsuccessful as the bolt and bushing continued to slip out of the arm on the motor shaft. Notified Scott Faas who sent some pictures of how the shaft should be attached. Removed the TDDR assembly to the lab to facilitate maintenance at 1500 GMT on 8/24. Removed the shaft assembly and compared with the pictures sent by Scott Faas. The pictures showed a flat washer which should prevent the bolt head from slipping through the bushing in the arm. This washer was not present on the GRAMS OT assembly. Furthermore, if a washer was installed the bolt head would protrude through the arm too far and not allow clearance for the arm to rotate. Scott advised that the bushing on the arm that he had was pressed tight into the arm. The bushing on GRAMS OT assembly slipped freely through the arm. Cleaned the bushing and arm with alcohol and used Loctite to assemble the two pieces together. Also cleaned and oiled the shaft bearing to free it up. Reassembled the unit and placed back into GRAMS OT at 2000 GMT. The GRAMS program was not running at that time due to high temperature in the enclosure. RTS at 2015 GMT. Will monitor shadow band to ensure that it rotates properly when the GRAMS software is running.

8/19/99

CM

T/S shadow band problem and found that the arm which connects between the motor shaft and shadow band gears had come loose at the motor end. Reconnected the arm but it came loose again after a few minutes. The arm will not stay attached to the motor. Contacted Scott Faas who will try to locate a spare arm assembly for replacement or determine if the current assembly can be fixed. Shadow band remains inop at this time.

8/19/99

ZW

The TDDR Shadow band was not moving when it was checked on morning rounds. Left voice mail with Scott Faas. RTS 1530 GMT.

8/6/99

MR

Performed the weekly and monthly PM.

6/18/99

MR

Performed weekly PM. Pump pressure is low but seems to be working right. RTS at 1350.

6/14/99

CAM

SDS comlog reports indicated that GRAMS OT had not collected since 1301 GMT on 6/12. Removed the GRAMS enclosure cover and found that the program was locked up. All channels were reading 0 and not changing. Rebooted the computer at 1325 GMT. RTS status O.K. at 1330 GMT.

5/14/99

JEM

Performed weekly PM. RTS @1353 GMT.

1/26/99

MR

PM. The air pump stopped working again. Reset the circuit breaker and the pump started but stopped shortly after. R/R the *pump. RTS at 2100. *Old - 9708008450 *New - 9708008449.

1/25/99

MR

PM. The air pump has stopped working. Reset the circuit breaker and the pump started to work.

1/4/99

MR

The grams pump has seized up. R/R the pump. Removed and cleaned the cooling fan screens and R/R one of the cooling fans.

12/3/98

KT

At 2100 GMT the software had locked up on the instrument preventing any data to get collected. At 2105 GMT the system was rebooted and returned to service. RTS at 2111 GMT.

10/14/98

SF/MF

Scott Faas and Mike Ferrario on site and report the following: Booster fan to increase cooling flow to air conditioner installed. Insulation below radiometer deck installed. *TDDR installed. *New - 09

9/23/98

KT

At 1358 GMT re-connected a piece of insulation that had blown off of the instrument housing.

9/11/98

SS

Weekly maintenance revealed that the pump pressure was low. Found the out filter to be leaking air around the seal. Turned off the pump tightened the filter and restored power to the pump. Pump pressure now reads 8 PSI. RTS at 1445 GMT.

8/28/98

SS

Weekly checks revealed that the pump was not working properly. Found the pump outlet filter plug was broken out of the pump. Removed the remaining potion of the plug and replaced it with a spare. RTS at 1515 GMT.

8/28/98

CM/GS

Instrument mentor requested that the GRAMS OT air conditioner be removed and sent in to local heat and air firm for service to check operation and refrigerant charge. Air conditioner was removed on 8/26/98 at 1603 GMT. Received the air conditioner back on 8/28/98. The service technician advised that they could find nothing wrong with the air conditioner and the refrigerant was fully charged. At mentors request the pressure switch was changed from 24 to 30. Installed GRAMS OT air conditioner from 1730 to 1845 GMT. The insulation that was previously placed on the side of the enclosure was removed but the thin reflective insulation blanket was left in place on the deck. Powered the system up at 1845 GMT. At that time the enclosure temperature was 94F so the UPS had cut AC power. Remote status check at 2105 GMT found the system was then operational.

8/26/98

CAM/MW

GRAMS OT has consistently shut down during the heat of the afternoon due to the air conditioner being unable to keep the enclosure below 85F.Instrument mentor requested that the GRAMS air conditioner be removed and taken in for service. The compressor will be checked out and the refrigerant charged if necessary. In addition the pressure switch will be changed from 24 to 30. Removed GRAMS OT enclosure covers at 1600 GMT. Powered the system down at 1603 GMT. Removed the air conditioner to be sent in for service. Replaced the covers at 1620 GMT and taped plastic over the open hole where the air conditioner inlet was originally. System off line until air conditioner is returned.

8/25/98

SS

Removed the air conditioner inlet and exhaust filters at mentors request.

8/20/98

CAM

Instrument mentor requested that an insulating layer be placed over the deck of GRAMS OT to further reduce the effect of external heat loading on the air conditioner. The system has consistently shut down in the afternoons due to high enclosure temperature. Covered the GRAMS OT deck with foil lined bubble insulation from 1545 to 1605 GMT. RTS at 1605 GMT.

8/18/98

CAM

GRAMS OT continues to shut down for a couple of hours each afternoon due to high enclosure temperature. Instrument mentor requested that insulation be placed on the south side of the GRAMS enclosure to reduce external heat loading. Attached 2" thick poly foam to the east and south sides of the enclosure and wrapped the whole enclosure in foil coated bubble film to further reduce direct solar heating. Maintenance performed between 1430 and 1500 GMT. Will continue to monitor status.

8/14/98

CAM

GRAMS OT collection failed at 1901 GMT on 8/14/98. The system had shut down due to high internal temperature. Measured the enclosure temperature using a calibrated hand held sensor and found that the Athena temperature controller was reading approximately 8 degrees high. Replaced the controller. New controller agrees with hand held sensor however the enclosure temperature is still running higher than it should. Bypassed the air conditioner external controls and configured the unit so that the compressor should run continuously. The compressor would not run continuously but rather cycled on and off in very short intervals. Returned the unit to its original configuration and RTS at 2100 GMT. T/S will continue.

8/13/98

CAM

GRAMS OT has been shutting for a few hours each afternoon down due to high enclosure temperature since 7/18/98. Powered GRAMS OT down at 1740 GMT on 8/13/98. Removed the air conditioner for maintenance. Removed the cover and checked condenser and evaporator for dirt and obstructions. The air conditioner interior was observed to be clean with no obstructions. Checked the inlet/outlet temperature differential to be approximately 25 degrees F. The outlet temperature was 52F with an inlet temperature of 77F. Installed the air conditioner back in the enclosure and used Styrofoam and tape to help provide a more efficient air flow pattern in the enclosure. When the system was powered back on the Athena Temperature controller power was intermittent. Tightened two of the contacts on the controller sleeve and secured the controller din mount more firmly. System RTS at 2200 GMT. Will continue to monitor status.

8/13/98

CAM

GRAMS OT has been shutting for a few hours each afternoon down due to high enclosure temperature since 7/18/98. Powered GRAMS OT down at 1740 GMT on 8/13/98. Removed the air conditioner for maintenance. Removed the cover and checked condenser and evaporator for dirt and obstructions. The air conditioner interior was observed to be clean with no obstructions. Checked the inlet/outlet temperature differential to be approximately 25 degrees F. The outlet temperature was 52F with an inlet temperature of 77F. Installed the air conditioner back in the enclosure and used styrofoam and tape to help provide a more efficient air flow pattern in the enclosure. When the system was powered back on the Athena Temperature controller power was intermittent. Tightened two of the contacts on the controller sleeve and secured the controller din mount more firmly. System RTS at 2200 GMT. Will continue to monitor status.

7/28/98

CAM

SDS comlog indicated that no files had been received from GRAMS OT since 0801 GMT on 7/28/98. The error 'Collection Failed - No Files to Collect' has occurred on each collection period since then. Remote status check via SMS software found that GRAMS OT program was hung up. Toggled the restart switch at 1312 GMT. RTS at 1315 GMT. Remotely checked status again at 1400 GMT and discovered that the TSBR and FSBR TTL's did not seem to be activated. Completely shut down the program and restarted at 1408 GMT. TTL's then went active and everything looked O.K. RTS at 1410 GMT.

7/27/98

CAM

GRAMS OT status check via SDS comlog revealed that no files had been collected since 0301 GMT on 7/27/98. The error 'Collection Failed - No Files to Collect' occurred. Checked instrument program status remotely using SMS software. Found that the programs run/shutdown switch was in the shutdown position. Returned the switch to the run position at 1335 GMT. Restarted GRAMS software and RTS status O.K. at 1345 GMT.

6/26/98

SS

PM visual revealed that the pump pressure was low. Inspected pump no problem found. Adjusted the pump output pressure to 8.5.

6/29/98

CAM

SDS comlog indicated that no files had been collected from GRAMS OT since 0303 GMT on 6/29/98. Remote check of instrument status using SMS software found that the run/shutdown switch was in the shutdown position. Moved the switch to the run position but there was still no data being generated and the system appeared to be hung up. Shut down the program and restarted software at 1350 GMT. System RTS status O.K. at 1353 GMT.

6/26/98

CAM

The GRAMS OT Athena temperature controller was installed in GRAMS RCF on 6/25 as part of T/S to isolate cause of erroneous temperature readings. Installed the original GRAMS OT controller at 1400 GMT on 6/26.

6/25/98

CAM

The GRAMS RCF has been shutting down during mid afternoons with a high temperature alarm. The internal process temperature reports 20-25 degrees higher than the actual internal temperature as checked with a calibrated handheld. Initial troubleshooting points towards a failed Athena controller or possibly the temperature sensor. In an effort to further isolate the failed component, the Athena controller was removed from the GRAMS RCF at 1908 GMT and installed in GRAMS OT at 1930 GMT. The GRAMS OT Athena controller was then installed in the GRAMS RCF at 1938 GMT. Preliminary indications now are that the GRAMS OT is showing the same characteristics as GRAMS RCF was originally and GRAMS RCF is operational. GRAMS OT has now shut down due to high temperature alarm. At this time it appears that it was the controller that was causing the erroneous temperature readings. Will allow systems to stabilize and continue to monitor both with current set up for 24 hours and order replacement parts as necessary.

6/25/98

CAM

SDS comlog revealed that no files had been collected from GRAMS OT since 0201 GMT on 6/25/98. Opened GRAMS OT enclosure at 1357 GMT. Found that the run/shutdown switch on the GRAMS25.VI interface was in the shutdown position. Moved switch to the run position at 1400 GMT. Also installed SMS software on the GRAMS PC to allow remote monitoring and control capabilities. Sealed enclosure and RTS at 1420 GMT. Check of SDS comlog indicated that a file was successfully collected during the next collection after returning the switch to run mode at 1401 GMT.

5/15/98

SS

Due to power failure power cycled instrument at 1634 GMT.

1/26/98

SS/GS

Observed level bubble out of inner circle but still within outer circle. leveled unit bubble to within the inner circle.

1/16/98

KT

Performed weekly on the GRAMS at 1642 GMT.

11/19/97

Mentor

Ventilator housing and pump reinstalled by mentor and fully operational.

11/6/97

CAM

There is light condensation inside the TDDR dome. This is an ongoing occurrence and mentor is aware of the situation.

10/31/97

CAM

Performed the weekly PM at 1510 GMT. Status O.K.

10/28/97

CAM

The TDDR dome has light condensation inside. Notified mentor of the situation. Mentor advised that according to Scripps comments regarding the flying versions of these radiometers, a little condensation will not hurt the measurements much. However, we do need to keep an eye on the situation and purge the head if it gets worse. No corrective action taken at this time.

10/23/97

TMG

Changed network address from 198.124.98.65 to 198.124.99.55. Installed SMS software for remote monitoring. Work performed on Grams-Cf from 1515 GMT to 1545 GMT resulted in data loss from rebooting the computer several times. RTS at 1545 GMT.

10/10/97

CAM

Performed the weekly and monthly PM checks at 1630 GMT. Cleaned air conditioner filters. Status O.K.

9/12/97

MW

The GRAMS OT shadow band is not rotating properly. The band is positioned horizontally to the West and is attempting to rotate but it only moves approximately 1 centimeter towards the East then returns to its original position. Powered the system off at 1626 GMT.

9/8/97

MW

The air compressor is inoperable and one of its associated cooling fans is making an unusual noise. The compressor feels very warm to the touch. Turned the GRAMS circuit breaker off at 2100 GMT. No further corrective maintenance performed at this time.